# thebibliography

### v3

@@ III-A Specifications
We first provide details of the specifications we have used in our QNNs design to carry out the analysis.
@@ VI Conclusion
This paper presents a significant contribution to the understanding and enhancement of Quantum Neural Networks (QNNs) training in the presence of quantum noise, a prevalent challenge in the Noisy Intermediate Scale Quantum (NISQ) era. Our findings underscore the pivotal role of quantum noise in precipitating the onset of barren plateaus (BPs), which are known to hinder the scalability and trainability of QNNs. Key insights from our research include the finding that the use of the PauliZ observable in a global cost function setting (all qubit measurement) maintains satisfactory trainability up to 6 qubits, and moderate trainability at 8 qubits, under noisy conditions. However, it also succumbs to BPs at 10 qubits. In contrast, the PauliX and PauliY observables, regardless of the cost function definition, exhibit substantially flatter optimization landscapes with increasing qubit numbers, indicating a quicker onset of BPs in noisy environments.
Our most notable finding is the advantageous use of a customized Hermitian observable combined with a global cost function. This approach not only mitigates the detrimental effects of quantum noise but also exploits these noisy conditions to enhance QNN trainability, effectively up to the 10-qubit limit of our study. Furthermore, in a local cost function setting, the PauliZ observable outperforms others, maintaining efficient training capabilities up to 10 qubits.
These results highlight the importance of careful observable selection and cost function definition in QNN training. Our work contributes to the advancement of quantum machine learning research, offering a path forward in optimizing QNNs for real-world applications in the NISQ era. We mainly highlight the importance of qubit measurement observable from the aspect of tacking the amplitude damping quantum noise. In near future, we aim to analyze more quantum noise to for a wide range of different target problems.

### v4

@@ VI Conclusion
This paper explores the impact of quantum noise on QNN trainability, highlighting how observable selection and cost function design influence the onset of barren plateaus (BPs) in the NISQ era. We show that PauliZ with a global cost function maintains good trainability up to 68686–86 – 8 qubits but fails at $10$, while PauliX and PauliY lead to earlier BPs. Our key finding is that a custom Hermitian observable with global cost not only resists noise-induced degradation but leverages noise to enhance training up to $10$ qubits. PauliZ also performs best under local cost settings. Future work will extend to other noise types and broader problem domains.